OpenStudy (anonymous):

During an intramural basketball game, Team A scored 12 fewer points than Team B. Together, both teams scored a total of 150 points. How many points did Team A score during the game?

OpenStudy (anonymous):

Ok, need help? I will help you :)

OpenStudy (anonymous):

look, Team A ( it will be just A or B the next time you see it) scored = B -12 overal it is A+B= 150, but because A is B-12. we get B-15 + B = 150 2B - 12 = 150 2B = 162 B= 81 A = B - 12 A = 81 - 12 A = 69 i guess

OpenStudy (anonymous):

you got it?

OpenStudy (anonymous):

thanks alot

OpenStudy (anonymous):

You are welcome :)